Solve the equation:
5/x-3=4/2x-7

Answer by stanbon(75887)

5/(x-3) = 4/(2x-7)
Cross-multiply to get:
10x-35 = 4x-12
6x = 23
x = 23/6
================
Cheers,
Stan H.